# SlowFast-SCI: Slow-Fast Deep Unfolding Learning for Spectral Compressive Imaging

### v1

@@ Abstract
Humans learn in two complementary ways: a slow, cumulative process that builds broad, general knowledge, and a fast, on-the-fly process that captures specific experiences. Existing deep-unfolding methods for spectral compressive imaging (SCI) mirror only the slow component—relying on heavy pre-training with many unfolding stages—yet they lack the rapid adaptation needed to handle new optical configurations. As a result, they falter on out-of-distribution cameras, especially in bespoke spectral setups unseen during training. This depth also incurs heavy computation and slow inference. To bridge this gap, we introduce SlowFast-SCI, a dual-speed framework seamlessly integrated into any deep unfolding network beyond SCI systems. During slow learning, we pre-train or reuse a priors-based backbone and distill it via imaging guidance into a compact fast-unfolding model. In the fast learning stage, lightweight adaptation modules are embedded within each block and trained self-supervised at test time via a dual-domain loss—without retraining the backbone. To the best of our knowledge, SlowFast-SCI is the first test‐time adaptation–driven deep unfolding framework for efficient, self‐adaptive spectral reconstruction. Its dual‐stage design unites offline robustness with on‐the‐fly per‐sample calibration—yielding over 70% reduction in parameters and FLOPs, up to 5.79 dB PSNR improvement on out‐of‐distribution data, preserved cross‐domain adaptability, and a 4× faster adaptation speed. In addition, its modularity integrates with any deep-unfolding network, paving the way for self-adaptive, field-deployable imaging and expanded computational imaging modalities. Code and models are available at https://github.com/XuanLu11/SlowFast-SCI
@@ 1 Introduction
Humans are endowed with a complementary learning system mcclelland1995there ; hong2024slowfast that operates on two distinct timescales. A slow process in the neocortex gradually integrates diverse experiences into a coherent world model, enabling robust prediction of action outcomes schwesinger55review . In parallel, a fast process in the hippocampus rapidly encodes episodic details, supporting agile adaptation to novel circumstances tulving1983elements . Integrating such a dual-speed mechanism into neural networks enables rapid adaptation, i.e., swift recalibration to new domains or unseen inputs—thereby enhancing both flexibility and generalization.
@@
This dual-speed learning paradigm inspires advances in spectral compressive imaging (SCI), a computational photography technique that aims to reconstruct rich hyperspectral information from compressed measurements. The predominant SCI architecture, coded aperture snapshot spectral imaging (CASSI) gehm2007single_cassi ; wagadarikar2008single_cassi_sparsity ; meng2020endtsa_cassi , employs a spatially coded mask and a dispersive element to modulate, shift, and integrates spectral bands onto a two-dimensional sensor zhang2022herosnet , achieving high-throughput, cost-effective, and bandwidth-efficient acquisition. The resulting hyperspectral images (HSIs) deliver rich spatial–spectral signatures that significantly enhance scene characterization beyond conventional RGB imagery. Consequently, SCI has found broad applications across computer vision and remote sensing, e.g., object detection van2010tracking_object ; kim20123d_object , medical imaging lu2014medical ; Meng_2020_OL_SMEM_medical , and land-cover classification borengasser2007hyperspectral_remote ; yuan2017hyperspectral_remote .
@@
Based on the CASSI acquisition model, a broad spectrum of reconstruction algorithms has been proposed to recover a three-dimensional hyperspectral data cube from its two-dimensional measurement. Traditional model-based methods employ handcrafted priors—such as sparsity wagadarikar2008single_cassi_sparsity ; kittle2010multiframe_sparsity , total variation rudin1992nonlinear_tv ; wang2015dual_tv , and low-rank constraints liu2018rank_lowrank ; zhang2019computational_lowrank —but they require careful parameter tuning and often yield suboptimal quality and slow convergence. The advent of deep learning introduced end-to-end networks that significantly improve both reconstruction accuracy and speed mst ; cst ; dauhst ; hdnet ; meng2020gap ; meng2020endtsa_cassi ; bisci ; Dong_2023_CVPR_rdluf ; Zhang_2024_CVPR_ssr ; Zhang_2024_CVPR_dpu .
@@
In particular, deep unfolding frameworks (DUNs) unroll iterative optimization into trainable layers, thereby integrating model-driven regularization with data-driven learning meng2020gap ; huang2021deep ; Zhang_2024_CVPR_dpu ; dauhst ; Dong_2023_CVPR_rdluf . However, these methods depend on extensive pre-training over large, paired datasets—an impractical requirement given the scarcity and acquisition cost of real HSI data, reflecting only the slow component of learning. Moreover, to achieve state-of-the-art performance, deep unfolding approaches often employ a large number of sequential unfolding blocks; each block must process the full high-dimensional hyperspectral volume, which greatly increases computational and memory burdens. This level of complexity not only complicates training and inference but also makes such architectures impractical for real-world applications and deployment on resource-constrained or edge devices.
@@
Per-sample adaptation—analogous to human fast learning—is essential for robust SCI reconstruction under scarce HSI data. While recent advances such as self-supervised fine-tuning and test-time adaptation enable rapid calibration, applying them directly to deep unfolding networks proves problematic: full-network fine-tuning destabilizes sequential stages and linear probes cannot accommodate layer-wise inference. SAH-SCI liu2024sahsci supports general self-supervised tuning but lacks runtime adaptation, and recent test-time adaptation (TTA) methods zhao2024adapnet ; Qin_2023_CVPR ; quan2022dual have yet to address hyperspectral SCI.
@@
To overcome these challenges, we introduce SlowFast-SCI, the first test-time adaptation-driven deep unfolding framework for efficient, self-adaptive spectral reconstruction. In the slow learning stage, we pre-train a deep unfolding backbone that fuses model-based priors with learnable parameters to acquire robust, general reconstruction capabilities. In the fast learning stage, lightweight Fast TTA modules (FAST-TTAMs) are inserted into each unfolding block and self-supervisedly fine-tuned at test time, instantly aligning the model to new measurements while preserving the frozen backbone. By uniting offline robustness with on-the-fly adaptation, SlowFast-SCI achieves superior cross-domain generalization without additional labels.
Furthermore, DUNs typically rely on numerous sequential stages to ensure high reconstruction fidelity, but this depth comes at the cost of heavy computation and slow inference. To address this, we apply an imaging mechanism-guided model distillation step that compresses the pre‐trained backbone into a streamlined “fast unfolding” variant with significantly reduced complexity. We then integrate the same lightweight adaptation modules into this distilled network and optimize only these modules during inference using a dual‐domain self‐supervised loss quan2022dual . With the backbone frozen, this design delivers swift test‐time tuning and efficient inference, mirroring the human ability to rapidly leverage distilled knowledge for new experiences. We summarize our contributions as follows:
@@
  • We develop a lightweight self-supervised adaptation module embedded in each unfolding stage to enable rapid per-sample calibration at inference.
  • A model distillation pipeline that compresses the pre-trained backbone into a fast unfolding variant, reducing parameter count and FLOPs by over 70% while preserving adaptability.
  • Evaluation on standard Harvard, ICVL and real CASSI measurements, showing up to 5.79 dB PSNR gain on out‐of‐distribution setups and a $4\times$ speedup in adaptation time.
@@ 2 Related Work
Hyperspectral Image Reconstruction. Related work in hyperspectral image reconstruction comprises two main streams. Traditional optimization‐based methods employ hand‐crafted priors—such as sparsity wagadarikar2008single_cassi_sparsity ; kittle2010multiframe_sparsity , total variation rudin1992nonlinear_tv ; wang2015dual_tv , and low‐rank constraints liu2018rank_lowrank ; zhang2019computational_lowrank —to iteratively solve the inverse problem; while interpretable, they are limited by prior expressiveness and suffer from slow convergence. More recently, deep learning and hybrid approaches have accelerated SCI reconstruction mst ; cst ; dauhst ; hdnet ; Dong_2023_CVPR_rdluf ; Zhang_2024_CVPR_dpu , including end‐to‐end networks that map 2D measurements directly to hyperspectral volumes, plug‐and‐play frameworks that insert learned denoisers into iterative solvers, and deep unfolding architectures—e.g., DAUHST dauhst , RDLUF Dong_2023_CVPR_rdluf , and DPU Zhang_2024_CVPR_dpu —that unroll optimization iterations into trainable layers. These learning‐based methods deliver state‐of‐the‐art accuracy and speed but depend heavily on large paired datasets of hyperspectral images and measurements, making them brittle under optical variations or spectral distribution shifts at test time. In contrast, our work introduces a complementary paradigm: a self‐supervised, test‐time adaptation framework that enables rapid per‐sample calibration without additional labeled data, thereby improving robustness and generalization across unseen scenes.
@@
Test-Time Adaptation. TTA aims to adapt a pre-trained model during inference to better align it with test data, improving performance without access to labeled test sample sun2020test_tta ; darestani2022test_tta . In image reconstruction, TTA methods have been introduced to fine-tune pre-trained models at test time quan2022dual ; Qin_2023_CVPR ; zhao2024adapnet . AdaptNet zhao2024adapnet embeds self-supervised fine-tuning blocks into deep unfolding networks, yielding strong performance in CS-MRI and SV-CT, but its considerable depth leads to slow adaptation. In contrast to previous inefficient approaches, we propose a lightweight adaptation module and a concise distillation pipeline, which together enable efficient test-time adaptation while exploiting the strengths of adapter architectures.
@@
Model Distillation. Model distillation buciluǎ2006model_distill ; hinton2015distilling is a widely adopted technique for compressing large models into smaller ones by transferring knowledge from a pre-trained teacher to a student network. While early methods focused on soft-label supervision for classification tasks, subsequent work has extended distillation to various forms, including intermediate feature matching romero2014fitnets , attention transfer Zagoruyko2017AT , and data-free distillation zhu2021data . Distillation has also been applied to self-supervised learning caron2021emerging , multi-task learning li2020knowledge , and test-time adaptation wang2021tent . Existing distillation techniques primarily target semantic or classification tasks, leaving inverse problems—especially SCI‐based image reconstruction—largely unexplored. We address this by introducing a structural‐knowledge distillation framework that transfers imaging priors from a pre-trained backbone into a lightweight, fast‐unfolding student model. As a result, the student inherits the teacher’s robustness while remaining both efficient and adaptable during test‐time inference.
Figure 2: Illustration of SlowFast-SCI. The left side depicts the slow learning process, pre-training on the source data with a deep unfolding network. The right side shows the fast learning process, where FAST-TTAM rapidly adapts the fast unfolding network to align the model with target scenes at inference time. The far right illustrates the details of FAST-TTAM, which takes both spatial and spectral information into account during fast learning.
3 Method
SlowFast-SCI tackles two complementary challenges in SCI reconstruction with dual-speed learning. In the slow learning stage (Sec. 3.1), we overcome the ill-posed inverse problem and scarce HSI training data by pre-training a physics-guided unfolding backbone that fuses model priors with learnable parameters to learn robust, general reconstruction capabilities. In the fast learning stage (Sec. 3.2), we address domain shifts and per-sample variability by inserting lightweight spectral adapters into each unfolding block and self-supervisedly fine-tuning only these adapters at test time—instantly calibrating to unseen measurements while keeping the backbone frozen. Furthermore, we propose an imaging-guided distillation framework that condenses the hybrid architecture into a compact “fast unfolding” network, augmented with tailored spectral adaptation blocks. This design enables efficient, label-free reconstruction across domains, mirroring the rapid application of distilled expertise in human learning.
3.1 Slow Learning
Figure 3: The schematic of CASSI system.
@@ 3.1 Slow Learning
Mathematical Model of CASSI. SCI is a snapshot modality that captures a full hyperspectral data cube in a single coded measurement by jointly modulating and integrating across all spectral bands. As illustrated in Fig. 3, the CASSI implements SCI as follows. Let a hyperspectral patch consist of $N_{\lambda}$ bands $\{X_{i}\}_{i=1}^{N_{\lambda}}$, each $X_{i}\in\mathbb{R}^{H\times W}$. Each band is first modulated by a known binary mask $M$, then the modulated frames are shifted along the detector by d(i1)d(i-1) pixels and summed to yield the compressed measurement $Y(u,v)=\sum_{i=1}^{N_{\lambda}}M\odot X_{i}\bigl(u,\,v+d(i-1)\bigr)+N(u,v),$ where $\odot$ denotes the Hadamard product, $N$ is additive sensor noise, and $(u,v)$ index spatial coordinates. By vectorizing the 3D hyperspectral cube into $x\in\mathbb{R}^{nN_{\lambda}}$ and the 2D measurement into $y\in\mathbb{R}^{n}$ with $n=H\bigl(W+d\,(N_{\lambda}-1)\bigr),$ the acquisition model can be written in matrix–vector form as
@@
Deep Unfolding Network-driven Slow Learning. Traditional optimization-based spectral reconstruction guarantees convergence and interpretability but tends to be brittle; CNNs and transformers add flexibility but require large labeled datasets—unrealistic in spectral imaging. Deep unfolding unrolls a MAP optimizer into a trainable network, marrying solver stability with learned representational power to achieve high-fidelity reconstructions from limited data. Therefore, we adopt this MAP-based unfolding as our slow-learning backbone for rapid adaptation.
@@
$\hat{\mathrm{\mathbf{x}}}_{\mathit{map}}=\arg\min_{\mathrm{\mathbf{x}}}\log P(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})=\arg\min_{\mathrm{\mathbf{x}}}\log P(\mathrm{\mathbf{y}}|\mathrm{\mathbf{x}})+\log P(\mathrm{\mathbf{x}}).$ (2)

### v2

@@ Abstract
Humans learn in two complementary ways: a slow, cumulative process that builds broad, general knowledge, and a fast, on-the-fly process that captures specific experiences. Existing deep-unfolding methods for spectral compressive imaging (SCI) mirror only the slow component—relying on heavy pre-training with many unfolding stages—yet they lack the rapid adaptation needed to handle new optical configurations. As a result, they falter on out-of-distribution cameras, especially in bespoke spectral setups unseen during training. This depth also incurs heavy computation and slow inference. To bridge this gap, we introduce SlowFast-SCI, a dual-speed framework seamlessly integrated into any deep unfolding network beyond SCI systems. During slow learning, we pre-train or reuse a priors-based backbone and distill it via imaging guidance into a compact fast-unfolding model. In the fast learning stage, lightweight adaptation modules are embedded within each block and fine-turned self-supervised at test time via a self-supervised loss—without retraining the backbone. To the best of our knowledge, SlowFast-SCI is the first test‐time adaptation–driven deep unfolding framework for efficient, self‐adaptive spectral reconstruction. Its dual‐stage design unites offline robustness with on‐the‐fly per‐sample calibration—yielding over 70% reduction in parameters and FLOPs, up to 5.79 dB PSNR improvement on out‐of‐distribution data, preserved cross‐domain adaptability, and a 4× faster adaptation speed. In addition, its modularity integrates with any deep-unfolding network, paving the way for self-adaptive, field-deployable imaging and expanded computational imaging modalities. The models, datasets, and code are available at https://github.com/XuanLu11/SlowFast-SCI.
@@ 1 Introduction
Humans are endowed with a complementary learning system [22, 15] that operates on two distinct timescales. A slow process in the neocortex gradually integrates diverse experiences into a coherent world model, enabling robust prediction of action outcomes [32]. In parallel, a fast process in the hippocampus rapidly encodes episodic details, supporting agile adaptation to novel circumstances [34]. Integrating such a dual-speed mechanism into neural networks enables rapid adaptation, i.e., swift recalibration to new domains or unseen inputs—thereby enhancing both flexibility and generalization.
@@
This dual-speed learning paradigm inspires advances in spectral compressive imaging (SCI), a computational photography technique that aims to reconstruct rich hyperspectral information from compressed measurements. The predominant SCI architecture, coded aperture snapshot spectral imaging (CASSI) [13, 36, 24], employs a spatially coded mask and a dispersive element to modulate, shift, and integrates spectral bands onto a two-dimensional sensor [45], achieving high-throughput, cost-effective, and bandwidth-efficient acquisition. The resulting hyperspectral images (HSIs) deliver rich spatial–spectral signatures that significantly enhance scene characterization beyond conventional RGB imagery. Consequently, SCI has found broad applications across computer vision and remote sensing, e.g., object detection [35, 18], medical imaging [21, 25], and land-cover classification [2, 39].
Figure 2: Illustration of SlowFast-SCI. The left side depicts the slow learning process, pre-training on the source data with a deep unfolding network. IGFUD distills the knowledge of the deep unfolding backbone into a compact fast unfolding network (FastUN). The right side shows the fast learning process, where FAST-TTAM rapidly adapts the FastUN to align the model with target scenes at inference time. The far right illustrates the details of FAST-TTAM, which takes both spatial and spectral information into account during fast learning.
@@
Based on the CASSI acquisition model, a broad spectrum of reconstruction algorithms has been proposed to recover a three-dimensional hyperspectral data cube from its two-dimensional measurement. Traditional model-based methods employ handcrafted priors—such as sparsity [36, 19], total variation [31, 38], and low-rank constraints [20, 44]—but they require careful parameter tuning and often yield suboptimal quality and slow convergence. The advent of deep learning introduced end-to-end networks that significantly improve both reconstruction accuracy and speed [5, 4, 6, 16, 23, 24, 7, 12, 43, 42].
@@
In particular, deep unfolding frameworks (DUNs) unroll iterative optimization into trainable layers, thereby integrating model-driven regularization with data-driven learning [23, 17, 42, 6, 12]. However, these methods depend on extensive pre-training over large, paired datasets—an impractical requirement given the scarcity and acquisition cost of real HSI data, reflecting only the slow component of learning. Moreover, to achieve state-of-the-art performance, deep unfolding approaches often employ a large number of sequential unfolding blocks; each block must process the full high-dimensional hyperspectral volume, which greatly increases computational and memory burdens. This level of complexity not only complicates training and inference but also makes such architectures impractical for real-world applications and deployment on resource-constrained devices.
@@
Per-sample adaptation—analogous to human fast learning—is essential for robust SCI reconstruction under scarce HSI data. While recent advances such as self-supervised fine-tuning and test-time adaptation enable rapid calibration, applying them directly to deep unfolding networks proves problematic: full-network fine-tuning destabilizes sequential stages and linear probes cannot accommodate layer-wise inference. SAH-SCI [41] supports general self-supervised tuning but lacks runtime adaptation, and recent test-time adaptation (TTA) methods [47, 27, 28] have yet to address hyperspectral SCI and suffer from slow inference.
@@
To overcome these challenges, we introduce SlowFast-SCI, the first test-time adaptation-driven deep unfolding framework for efficient, self-adaptive spectral reconstruction. In the slow learning stage, we pre-train a deep unfolding backbone that fuses model-based priors with learnable parameters to acquire robust, general reconstruction capabilities. Furthermore, DUNs typically rely on numerous sequential stages to ensure high reconstruction fidelity, but this depth comes at the cost of heavy computation and slow inference. To address this, we apply an imaging mechanism-guided fast unfolding distillation step that compresses the pre‐trained backbone into a streamlined “fast unfolding” variant with significantly reduced complexity. In the fast learning stage, lightweight Fast TTA modules (FAST-TTAMs) are inserted into each fast unfolding block and self-supervisedly fine-tuned at test time, instantly aligning the model to new measurements while preserving the frozen backbone. By uniting offline robustness with on-the-fly adaptation, SlowFast-SCI achieves superior cross-domain generalization without additional labels.
With the backbone frozen, this design delivers swift test‐time tuning and efficient inference, mirroring the human ability to rapidly leverage distilled knowledge for new experiences. We summarize our contributions as follows:
@@
  • A imaging-guided fast unfolding distillation pipeline that compresses the pre-trained backbone into a fast unfolding variant, reducing parameter count and FLOPs by over 70% while preserving adaptability.
  • We develop a lightweight self-supervised adaptation module embedded in each fast unfolding stage to enable rapid per-sample calibration at inference.
  • Evaluation on standard Harvard, ICVL and real CASSI measurements, showing up to 5.79 dB PSNR gain on out‐of‐distribution (OOD) setups and a $4\times$ speedup in adaptation time.
@@ 2 Related Work
Hyperspectral Image Reconstruction. Related work comprises two main streams. Traditional optimization-based methods employ hand-crafted priors—such as sparsity [36, 19], total variation [31, 38], and low-rank constraints [20, 44]—to iteratively solve the inverse problem; while interpretable, they are limited by prior expressiveness and suffer from slow convergence. Deep learning and hybrid approaches accelerate SCI reconstruction [5, 4, 6, 16, 12, 42], including end-to-end networks, plug-and-play frameworks, and deep unfolding architectures (e.g., DAUHST [6], RDLUF [12], DPU [42]). Although these learning-based methods deliver state-of-the-art accuracy and speed, they rely heavily on large paired datasets, making them brittle under optical variations or spectral distribution shifts at test time. In contrast, our work introduces a complementary paradigm: a self-supervised, test-time adaptation framework that enables rapid per-sample calibration without additional labeled data, thus improving robustness across unseen scenes.
@@
Test-Time Adaptation. TTA adapts a pre-trained model during inference to improve performance without labeled test samples [33, 11]. In image reconstruction, TTA fine-tunes models at test time [28, 27, 47]. AdaptNet [47] integrates self-supervised test time training into deep unfolding networks, but its considerable depth leads to slow adaptation. In contrast, we propose a lightweight adaptation module and a concise imaging-guided fast unfolding distillation pipeline, enabling efficient test-time adaptation via adapter architectures.
@@
Model Distillation. Model distillation [3, 14] is a widely adopted method for transferring knowledge from a large teacher to a smaller student network. While initial methods used soft-label supervision, subsequent work extended distillation via intermediate feature matching [30] and attention transfer [40], and applied it to self-supervised learning [8] and TTA [37]. Existing distillation primarily targets semantic tasks, leaving inverse problems—especially SCI-based image reconstruction—largely unexplored. We address this by introducing a structural-knowledge distillation framework that transfers imaging priors from a pre-trained backbone into a lightweight, fast-unfolding student model. The student thus inherits the teacher’s robustness while remaining both efficient and adaptable during test-time.
3 Preliminaries
@@
Mathematical Model of CASSI. SCI is a snapshot modality that captures a full hyperspectral data cube in a single coded measurement by jointly modulating and integrating across all spectral bands. Let a hyperspectral patch consist of $N_{\lambda}$ bands $\{X_{i}\}_{i=1}^{N_{\lambda}}$, each $X_{i}\in\mathbb{R}^{H\times W}$. Each band is first modulated by a known binary mask $M$, then the modulated frames are shifted along the detector by d(i1)d(i-1) pixels and summed to yield the compressed measurement $Y(u,v)=\sum_{i=1}^{N_{\lambda}}M\odot X_{i}\bigl(u,\,v+d(i-1)\bigr)+N(u,v),$ where $\odot$ denotes the Hadamard product, $N$ is additive sensor noise, and $(u,v)$ index spatial coordinates. By vectorizing the 3D hyperspectral cube into $x\in\mathbb{R}^{nN_{\lambda}}$ and the 2D measurement into $y\in\mathbb{R}^{n}$ with $n=H\bigl(W+d\,(N_{\lambda}-1)\bigr),$ the acquisition model can be written in matrix–vector form as:
@@
$\hat{\mathrm{\mathbf{x}}}_{\mathit{map}}=\arg\min_{\mathrm{\mathbf{x}}}\log P(\mathrm{\mathbf{y}}|\mathrm{\mathbf{x}})+\log P(\mathrm{\mathbf{x}}).$ (2)
@@
4 Method
SlowFast-SCI tackles two complementary challenges in SCI reconstruction with dual-speed learning. In the slow learning stage (Sec. 4.1), we overcome the ill-posed inverse problem and scarce HSI training data by pre-training a physics-guided unfolding backbone that fuses model priors with learnable parameters to learn robust, general reconstruction capabilities. In the fast learning stage (Sec. 4.2), we address domain shifts and per-sample variability by inserting lightweight spectral adapters into each unfolding block and self-supervisedly fine-tuning only these adapters at test time—instantly calibrating to unseen measurements while keeping the backbone frozen. Furthermore, we propose an imaging-guided fast unfolding distillation framework that condenses the hybrid architecture into a compact “fast unfolding” network, augmented with tailored spectral adaptation blocks. This design enables efficient, label-free reconstruction across domains, mirroring the rapid application of distilled expertise in human learning.
Figure 3: Comparison of DUN methods on 10 scenes of Harvard dataset. For all methods, PSNR (dB) and per-scene inference time (ms) are reported. For the unfolding method DPU, results with different numbers of stages are provided.
4.1 Slow Learning
In the slow learning stage, we pre-train a physics-guided unfolding backbone that integrates analytical priors with learnable parameters to address the ill-posed nature of spectral reconstruction and the scarcity of hyperspectral data. By explicitly embedding the underlying imaging physics into the network design, the unfolding backbone acquires a robust and generalizable reconstruction capability even under limited supervision.
Table 1: Comparison between the end-to-end (E2E) network and the deep unfolding network (DUN), both employing the same Transformer backbone.
Method Category PSNR(dB)
MST [5] E2E 31.18
DAUHST [6] DUN 32.64
As illustrated in Tab. 1, when employing the same backbone architecture (e.g., Transformer), the DUN generally achieves superior performance compared to its end-to-end counterpart. Furthermore, DUNs operate in a multi-stage iterative manner (e.g., 3, 5, or 9 stages), where increasing the number of stages progressively refines reconstruction performance at the cost of higher inference latency, as shown in Fig. 3. Building upon this foundation, the slow learning stage yields a 9-stage deep unfolding backbone with strong reconstruction accuracy. However, directly deploying this full-depth model in the fast learning stage is suboptimal, as its considerable computational overhead limits efficiency and impedes rapid adaptation to novel scenes. To address these limitations, we propose Imaging-Guided Fast Unfolding Distillation (IGFUD), which bridges the gap between high-fidelity reconstruction and real-time adaptability by distilling the knowledge of the deep unfolding backbone into a compact fast unfolding network, thereby reducing model complexity while enabling unsupervised adaptation to unseen imaging conditions.
Figure 4: Log-scaled visualization of 1/p-values for different k-values. The p-values obtained by performing a permutation test ($n=1000$) on the kernel matrices derived from the Maximum Mean Discrepancy (MMD) [29] between CAVE and the other three datasets (KAIST, Harvard, and ICVL). Each image (originally of size 256 × 256 × 28) is processed using average pooling with different window sizes $k$. Green bold values indicate statistically significant differences ($p<0.05$).
Imaging-Guided Fast Unfolding Distillation. The fast unfolding distillation (FUD) process is shown in Fig. 2. During this process, the pre-trained model, with its parameters frozen, serves as the slow unfolding network (SlowUN), while the two-stage model with the same architecture as the pre-trained model is used as fast unfolding network (FastUN). The dataset used for FUD is a dataset that the SlowUN has not seen during supervised slow learning, as shown in Fig. 5. Firstly, mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times(W+d\times(N_{\lambda}-1)}$ are entered into the SlowUN to obtain the results $\{\mathrm{\mathbf{X}}_{S,i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$, which will be used as a label to supervise the training of FastUN. Then mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ are input into FastUN to obtain the results $\{\mathrm{\mathbf{X}}_{F,i}\}_{i=1}^{N_{d}}$. We update FastUN by minimizing the following loss:
$\mathscr{L}_{FUD}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\|\mathrm{\mathbf{X}}_{F,i}-\mathrm{\mathbf{X}}_{S,i}\|_{F}^{2}.$ (6)
The inference speed of the two-stage FastUN obtained by FUD is much faster than the original pre-trained model, and its performance on test data is comparable to that of the original pre-trained model, as shown in Fig. 3.
Figure 5: Evaluating the cross-domain generalization of classic DUN. Images with blue backgrounds represent the average log-spectrum of each dataset. The deep unfolding network is trained on the CAVE dataset during the slow learning stage. To evaluate cross-domain generalization, the model is tested on two out-of-distribution (OOD) datasets (Harvard and ICVL) and on real-world test samples, rather than on in-distribution (IND) datasets such as KAIST used in prior works. The average PSNR (dB) demonstrates that classic DUN performs well on IND test set but experiences a noticeable performance drop when applied to OOD test samples.